Randomized Algorithms - Presentation Review

Julian Dymacek LCSEE, West Virginia University, Morgantown, WV {ksmani@csee.wvu.edu}

March 6, 2012

1 Instructions

(a) Give a score between 1 and 10 for each question, unless it is a non-scoring question.

(b) Please answer questions professionally and sincerely.

2 Review

1. Is the lecturer prepared?

2. Did the lecturer communicate the material to you in advance?

3. Does the lecturer know the material he is presenting?

4. Did the instructor answer your questions to your satisfaction?

5. Did the instructor maintain a good pace?

6. Did you prepare and come for the lecture?

7. Did you understand the material communicated by the instructor?

8. What are the strengths of instructor? (No score)

9. What are the weaknesses of instructor? (No score)

10. Overall, how would you rate this lecture?

11. Overall, how would you rate this lecturer?